# TimeFlow: Longitudinal Brain Image Registration and Aging Progression Analysis

### v2

@@ Conclusion.
In this work, we introduced TimeFlow, a novel framework for longitudinal brain MR registration that addresses key challenges in the field, including the trade-off between accuracy and temporal consistency, limited sequential data, and the inability to predict future brain states. TimeFlow leverages a UNet-based architecture with temporal condition embeddings to ensure temporally continuous and spatially smooth deformation fields, eliminating the need for explicit smoothness constraints. Unlike existing methods that require full sequential data, TimeFlow learns directly from even just two images per patient, achieving both high registration accuracy and robust extrapolation to future timepoints. Our results show that TimeFlow outperforms baselines in both interpolation and extrapolation tasks and facilitates meaningful brain aging and neurodegenerative disease progression analysis. By bridging interpolation and future state prediction, TimeFlow facilitates clinically relevant analyses without extra anatomical annotations, representing a significant step toward data-driven models for tracking, predicting, and ultimately intervening in neurodegenerative diseases.
Acknowledgements
This work was supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWi) through the project NeuroTEMP, and by the Munich Center of Machine Learning (MCML). Jiazhen Pan is funded by the European Research Council (ERC) grant Deep4MI (No. 884622).
The authors declare that they have no competing interests.
Supplementary Material
7 Pairwise Registration
Pairwise Baseline Endpoint
[…]
rigid 11.0$\pm$4.5 16.6$\pm$2.2 - -
Seq-VXM 7.0$\pm$4.1 21.2$\pm$2.9 3.4$\pm$0.8 0.8$\pm$2.7
NODER 8.0$\pm$3.9 20.0$\pm$2.6 3.2$\pm$1.3 18.8$\pm$89.6
VXM 7.3$\pm$4.5 20.9$\pm$2.9 3.0$\pm$0.6 15.2$\pm$73.2
VXMdiff 6.7$\pm$3.7 21.6$\pm$2.9 4.3$\pm$1.3 33.9$\pm$175.8
TimeFlow 7.3$\pm$4.5 20.8$\pm$2.9 3.2$\pm$0.2 2.0$\pm$4.1
TimeFlowdiff 6.6$\pm$4.2 21.7$\pm$3.0 4.5$\pm$0.5 0.3$\pm$0.5
Table 3: Quantitative registration results for Pairwise Registration between baseline and endpoint. Accuracy (foreground): MAE, PSNR; Smoothness: SDlogJ, NDV.
8 Temporal Continuous Deformation Fields
Figure 6: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 7.
Figure 7: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 6.
Figure 8: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 9.
Figure 9: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 8.
Figure 10: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 11.
Figure 11: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 10.

### v3

@@ 7 Conclusion
We introduced TimeFlow, a novel framework for longitudinal brain MR registration, specifically designed to address persistent challenges such as limited sequential data, the accuracy-temporal consistency trade-off, and the inability to predict future brain states. TimeFlow leverages a U-Net-based architecture with temporal condition embeddings to ensure temporally continuous and spatially smooth deformation fields, eliminating the need for explicit smoothness penalties. Unlike existing methods that require full sequential data, TimeFlow learns effectively from only two scans per patient, achieving both high registration accuracy and robust extrapolation to future timepoints. Extensive evaluations demonstrate TimeFlow’s superior performance over existing baselines in both interpolation and extrapolation tasks, as well as its ability to facilitate clinically meaningful analyses of brain aging and neurodegenerative diseases. By seamlessly bridging retrospective interpolation and prospective extrapolation, TimeFlow enables annotation-free analyses, representing a significant step toward data-driven models for tracking and predicting neuroanatomical changes in neurodegenerative diseases.